# Emulating Full Client Participation: A Long-Term Client Selection Strategy for Federated Learning

### v1

@@ III-B2 Client Selection Across Multiple Rounds
to evaluate the frequency of a client $i$ being selected throughout the total $T$ rounds. Note that $\{x_{i,t}\}$ forms a stochastic process, and the expectation operation $\mathbb{E}$ is applied in Eq. (6). Given two clients $i,j\in\mathbb{N}$, we use $|p_{i}-p_{j}|$ to quantify the difference between their frequencies of being selected.
Building on $\text{Dist}_{i,j}(t)$ and $p_{i}$, we employ the $\epsilon\text{-}\delta$-IF framework utilized in [35, 36] to quantify the fairness constraint. We evaluate a client selection strategy at the end of the $T$-th communication round. Specifically, given $\epsilon,\delta\geq 0$, a client selection strategy is of individual fairness if for every pair of clients $i,j\in\mathbb{N}$ with $\text{Dist}_{i,j}(T)\leq\epsilon$, the difference in their frequencies of being selected, namely, $p_{i}$ and $p_{j}$, should not exceed $\delta$. The mathematical formulation of $\epsilon\text{-}\delta$-IF is provided in Definition 1, and the impact of parameter selection for $\epsilon$ and $\delta$ is discussed in Section V-C.
@@
Note that $\epsilon\text{-}\delta$-IF requires examining Eq. (7) for all pairs of clients, leading to significant computational overhead and rendering subsequent optimization challenging. To address this issue, we propose to determine a reference client $i^{\star}$ for each client $i$, which exhibits the largest difference in the frequency of being selected among clients whose gradient distances are less than $\epsilon$. That is,
@@
Then, given a client $i$ and its reference client $i^{\star}$, a client selection strategy is certified as $\epsilon\text{-}\delta$-IF if
@@ III-B3 The Optimization Function
III-C Problem Transformation Under Lyapunov Optimization
To solve P1 in Eq. (10), we leverage Lyapunov optimization, a technique from control theory used to analyze the stability of dynamic systems [37, 9]. The main idea behind Lyapunov optimization is to break down the long-term time-averaged constraints in the optimization problem into constraints that can be adhered to in each communication round. This allows us to optimize the objective function directly in each round without violating the constraints specific to that round. Furthermore, by ensuring that these constraints are adhered to in each round, we can automatically achieve compliance with the long-term time-averaged constraints at the end of the $T$ round. By leveraging Lyapunov optimization, the problem stated as P1 in Eq. (10) is ultimately converted into the problem P3 in Eq. (23), which can be efficiently solved online in each communication round. Details of problem transformation are described below.
Before employing Lyapunov optimization, we remove the absolute value sign from the constraint in Eq. (9) and rephrase it as two equivalent constraints,
@@ III-C Problem Transformation Under Lyapunov Optimization
Lyapunov optimization typically involves three steps [37]. First, a virtual queue is defined for each time-averaged constraint. This step converts the time-averaged constraints into stability constraints for virtual queues. Second, a Lyapunov function is defined based on virtual queues, and the Lyapunov drift is introduced. This step breaks down the time-averaged constraints into requirements for Lyapunov drift in each communication round. Finally, the Lyapunov drift is integrated into the optimization objective, allowing us to minimize both the Lyapunov drift and the optimization objective simultaneously. In the following, we illustrate how to transform problem P1 using these three steps.
First, we define two virtual queues, namely $Z_{i}(t)$ for the constraint in Eq. (11) and $Q_{i}(t)$ for the constraint in Eq. (12), for all clients $i\in\mathbb{N}$. Specifically, these queues are initialized as $Z_{i}(0)=0$ and $Q_{i}(0)=0$, and updated according to the following rule
@@
By the introduction of $Z_{i}(t)$ and $Q_{i}(t)$, we have Theorem 2, which converts the long-term constraints on $x_{i,t}$ in Eq. (11) and (12) into the stability constraints for $Z_{i}(t)$ and $Q_{i}(t)$. The proof of Theorem 2 is provided in the supplementary file.
@@
To make $Z_{i}(t)$ and $Q_{i}(t)$ remain stable, an intuitive approach is to bound the increase of these queues in each round so that they could not grow to infinity. Based on this intuition, in our work, we define a global queue $\Theta(t)$, which stores the state of $Z_{i}(t)$ and $Q_{i}(t)$ for all clients. That is,
@@
Second, following Lyapunov optimization, we define the Lyapunov function [37] as
@@
At last, to ensure that $\Delta\left(\Theta(t)\right)$ remains sufficiently small, a straightforward method to combine the optimization objection with the drift $\Delta\left(\Theta(t)\right)$, and minimize both simultaneously. Mathematically, in our work, the optimization problem in Eq. (10) is transformed into
@@
Based on Theorem 3, the problem P2 in Eq. (18) is transformed into
@@
where $Z_{i}(t)$ and $Q_{i}(t)$ are defined in Eq. (13), $m_{i,t}$ and $n_{i,t}$ are defined in Eq. (20), and $V$ is a hyperparameter that balances the optimization function and the fairness constraint. With an appropriate $V$, the solution in Eq. (23) can infinitely approximate the optimal solution in Eq. (18) [38]. In the following, we select clients by minimizing $G(\mathbb{S}^{t})$ in Eq. (23) in each round.
@@ IV-A The Client Selection Strategy
Specifically, a set function $g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ is submodular if for every $A\subseteq B\subseteq\mathbb{N}$ and $\𝐵i\in\mathbb{N}\backslash B$ it holds that $g(A\cup\{i\})-g(A)>g(B\cup\{i\})-g(B)$. One typical example of a submodular function is the facility location function [40]. Suppose we aim to select locations from a set of positions =$\mathbb{N}=\{1,\ldots,N\}$ to open facilities and to serve a collection of $K$ users. If a facility is located at position $j$, the service it provides to user $i$ is quantified by $M_{i,j}$. Each user is assumed to select the facility with the highest service, and the total service provided to all users is modeled by the set function
$f(\mathbb{S})=\sum_{i=1}^{m}\max_{j\in\mathbb{S}}M_{i,j},$ (24)
$\overline{G}(\mathbb{S}^{t})=G(\{e\})-G(\mathbb{S}^{t}\cup\{e\}),$ (25)
@@
5:     Determine the reference client $i^{\star}_{t}$ for $i\in\mathbb{P}_{k}$
@@ IV-B Convergence Analysis
We provide a theoretical analysis to demonstrate the convergence ability of the proposed client selection strategy. The analysis uses FedAvg as the aggregation method, and it can be extended to other federated optimization methods as well. To implement the theoretical analysis, we establish six assumptions regarding the local models and data distribution heterogeneity among clients.
First, based on the effective greedy selection strategy in Algorithm 1, we assume that the estimation error between the client subset and the full client set (Eq. (2)) is small and can be quantified by a variable $\rho$, as stated in Assumption 1. This assumption has also been studied in the context of data coreset selection [32] and client selection methods [27]. Note that $\rho$ is used as a measure to characterize the quality of the estimation, and the analysis holds for any $\rho<\infty$.
Assumption 1. At each communication round $t$, we assume that the gradient aggregated from the selected subset of clients can provide a good approximation of the gradient aggregated from the full set, i.e.,
@@
Based on these assumptions, the proposed client selection strategy is demonstrated to converge to the global optimal parameter $\bm{w}^{\star}$ at a rate of $\mathcal{O}(1/t)$ for heterogeneous data settings, as stated in Theorem 4. The proof is provided in the supplementary file. The convergence rate $\mathcal{O}(1/t)$ is the same as existing client selection methods [27, 18, 43].
@@
The term $\rho$ encodes the estimation error in Eq. (2). When more clients are selected in a communication round, the subset of selected clients can better approximate the data distribution of the full client set, then the term $\rho$ is smaller. Particularly, $\rho$ becomes zero when all clients are selected in each round, i.e., $K=N$. We also analyze the impact of varying $K$ empirically and propose an adaptive client selection method [44]. However, to reduce communication overhead, the number of selected clients is usually less than the total number of available clients, i.e., $K<N$. In this case, $\rho$ remains a non-vanishing term. In experiments, we observe that the proposed client selection strategy guides the system to converge along a trajectory similar to that of full client participation, but there exists a non-diminishing solution bias dependent on $\rho$. This observation is consistent with our theoretical analysis.
@@ V-A Experimental Settings
We consider the cross-device federated learning scenario where a large number of clients exist, each with limited computational power. In our experiments, we fix the number of clients at $N=100$ and employ FedAvg [1] as the aggregation method. The proposed client selection strategy is evaluated on two datasets, FMNIST [45] and CIFAR-10 [46]. Following the approach in [18], we opt for basic and small-scale models to accommodate the clients’ restricted computational resources. Specifically, for the FMNIST dataset, we utilize a multilayer perception (MLP) with two hidden layers as the global model. For CIFAR-10, our global model adopts a convolutional neural network (CNN) architecture consisting of three convolutional layers, with the outputs of these layers fed into an MLP layer. Further details on training and hyperparameter settings can be found in the supplementary file.
@@
We explore four data partitioning methods, including an independently and identically distributed (IID) approach and three heterogeneous approaches. In the IID approach, we randomly partition the dataset into $N$ parts, with each client assigned one part. We select $K=5$ clients in each round. Three heterogeneous data partitioning approaches are as follows.
@@
(i) 1 Shard Per Client (1SPC). Following the setting in [1], we divide the dataset into $N$ shards, ensuring that data within a shard shares the same label. We randomly assign a shard to each client so that a client has data with only one label. This is the most heterogeneous case. In this case, we select $K=10$ clients in each round.
@@
(ii) 2 Shards Per Client (2SPC). Similar to the 1SPC setting, the dataset is partitioned into $2N$ shards, where data within each shard shares the same label. Clients are randomly assigned two shards, where most clients have data with two labels (two shards with different labels), while a few clients have data with only one label (two shards with the same label). In this case, we set $K=5$ in each round.
@@
(iii) Dirichlet Distribution (Dir). We partition the dataset based on a Dirichlet distribution parameterized by a concentration variable $\alpha$, where a smaller value of $\alpha$ indicates higher heterogeneity. In our work, we set $\alpha=0.8$. Under this setting, clients have data with all labels, but the data size of each label varies among clients. We determine the data size of each label for each client following the analysis in [47, 18]. In this case, we choose $K=5$ clients in each round.
Figure 2: Test accuracy on FMNIST and CIFAR-10 under the IID scenario.
@@
Note that the 1SPC and 2SPC scenarios address label shift heterogeneity, while the Dir scenario accounts for heterogeneity in both labels and dataset size. Furthermore, apart from the assumption of a fixed number of clients ($K$) to be selected, we also explore the impact of varying $K$ and propose an adaptive method to determine its value. Details of the adaptive method can be found in [44].
@@
TABLE I: Time overhead analysis on FMNIST in the 1SPC scenario.
Method Random AFL Powerd DFairFed
Time (ms) For Client Selection Only 0.038 0.153 0.016 0.529
Time (ms) For Completing a Communication Round 2019 2044 2033 2472
We conduct the experiments five times using different randomized initial conditions and present the average results. In our experiments, we set $\delta=0.1$ , $\epsilon=0.3$, and $V=0.9$ for FMNIST, and $\delta=0.05$, $\epsilon=0.05$, $V=0.7$ for CIFAR-10. We compare the proposed LongFed with three baseline methods: (1) the random selection strategy (Random) [1]; (2) active client selection strategy (AFL), where clients with higher training losses have higher probabilities of being selected [16]; (3) power-of-choice strategy (Powerd), where clients are selected based on the power of $d$ choices load balancing strategy [17]. These methods are commonly used as baselines in [18, 27]. Additionally, we compare our method with the full client participation method (Full) with $K=100$.
@@ V-B Experimental Results
V-B1 Convergence Analysis
The experimental results under the IID scenario are presented in Fig. 2, and the results under three heterogeneous data partitioning methods are presented in Fig. 3. In Fig. 2 and Fig. 3, the $x$ axis represents the communication round, and the $y$ axis displays the test accuracy of the global model. For a clearer comparison, we apply the moving average rule [48] with a sliding window size of 10 to smooth the results.
Fig. 2 reveals that in the IID scenario, these methods achieve similar performance. For three heterogeneous scenarios on the FMNIST dataset, as shown in Fig. 3 (a), (b), and (c), respectively, the proposed method achieves performance that is the most similar to that with full client participation. Moreover, it exhibits faster convergence speed and superior test accuracy compared to prior works. Notably, there is an approximate $20\%$ improvement in the 1SPC scenario and an $8\%$ improvement in the 2SPC and Dir scenarios. Similar trends are also observed for the CIFAR-10 dataset.
Figure 4: Visualization of the selected clients on the FMNIST dataset under the 1SPC scenario. (a) is the t-SNE plot of client embeddings, where clients are organized into 10 clusters. (b)-(f) display the results of the selected clients using different client selection strategies, with the chosen clients marked by black stars.
Figure 5: Visualization of the selected clients on the FMNIST dataset under the 2SPC scenario. (a) shows the t-SNE plot of client embeddings, where the clustering pattern is not evident. In (b)-(f), circles are drawn around the selected clients, indicating that clients covered by a circle can be represented by the corresponding selected client.
V-B2 Time Overhead Analysis
We evaluate the time overhead of the proposed client selection strategy, and the results are presented in Table I. The time overhead analysis is conducted on the FMNIST dataset under the 1SPC scenario, with similar observations across other scenarios. In Table I, the first row focuses on the time solely for client selection within a communication round, while the second row denotes the overall time required for a complete communication round, including client selection, local updates, and global aggregation.
First, examining the time solely for client selection, our proposed method exhibits only a marginal increase (less than 0.4ms) compared to existing methods. This increment is acceptable in most real applications. In addition, the time required for client selection (about 0.5ms) is almost negligible compared to the total time required to complete a communication round (about 2000ms). This highlights that the majority of time is allocated to local updates within clients. Second, examining the time required to complete a round, our proposed strategy imposes only a modest additional computational time (about 400 ms) compared to existing approaches. This is primarily because evaluating the distance $\text{Dist}_{i,j}(t)$ in Eq. (5) is time-consuming, especially when the dimension of $\nabla f_{i}(\bm{w}^{t})$ is high. The time overhead can be reduced by employing a more efficient method for calculating distances in high-dimensional gradient space. Nevertheless, the marginal increase in time for the proposed strategy is justified when considering its superior performance in client selection.
V-B3 Visualization of Client Selection Strategy
We provide visualizations of the selected clients to offer an interpretable analysis of our proposed strategy. As shown in Fig. 4, in the 1SPC scenario on the FMNIST dataset, baseline methods typically select two or more clients from the same cluster. In contrast, the proposed LongFed, as shown in 4 (e), selects one client from each cluster, which effectively approximates the data distribution of the full client set. Additionally, as shown in Fig. 4 (f), it chooses a diverse subset of clients in the multi-round selection, adhering to Principle II.
The results in the 2SPC scenario are illustrated in Fig. 5, where the clustering pattern is not evident. For baseline methods, the circles around selected clients often overlap and cover only a minority of clients. However, in our proposed method, as shown in Fig. 5 (e), these circles cover the majority of clients. This suggests that the selected clients provide a better approximation of the full client set compared to the baseline approaches. Moreover, as shown in Fig. 5 (f), the strategy selects a diverse subset of clients across multiple rounds, satisfying Principle II. Visualization results of other scenarios are provided in the supplementary file.
V-C Individual Fairness Analysis

### v2

@@ III-B2 Formulation of Principle II
to evaluate the frequency of a client $i$ being selected in $T$ rounds. Note that $\{x_{i,t}\}$ forms a stochastic process, and the expectation operation $\mathbb{E}$ is applied in Eq. (6).
Building on $\text{Dist}_{i,j}(t)$ and $p_{i}$, we employ the $\epsilon\text{-}\delta$-IF framework utilized in [33, 34] to quantify the individual fairness constraint. We evaluate a client selection strategy at the end of the $T$-th round. Specifically, given $\epsilon,\delta\geq 0$, a client selection strategy is of individual fairness if for any clients $i,j\in\mathbb{N}$ with $\text{Dist}_{i,j}(T)\leq\epsilon$, the difference in their selection frequencies, namely, $p_{i}$ and $p_{j}$, should not exceed $\delta$. The mathematical formulation of $\epsilon\text{-}\delta$-IF is provided in Definition 1, and the impact of parameter selection for $\epsilon$ and $\delta$ is discussed in Section V-D.
@@
The $\epsilon\text{-}\delta$-IF requires examining Eq. (7) for all pairs of clients, leading to significant computational overhead. To address this issue, we propose to determine a reference client $i^{\star}$ for each client $i$, which exhibits the largest difference in the selection frequency and has the gradient distance less than $\epsilon$. That is,
@@
Then, the $\epsilon\text{-}\delta$-IF is simplified as
@@ III-B3 The Optimization Function
III-C Transformation Under Lyapunov Optimization
To solve P1 in Eq. (10), we leverage Lyapunov optimization, a technique from control theory used to analyze the stability of dynamic systems [35, 8]. The main idea behind Lyapunov optimization is to break down the long-term time-averaged constraints into constraints that can be adhered to in each communication round. By leveraging Lyapunov optimization, the problem stated as P1 in Eq. (10) is ultimately converted into the problem P3 in Eq. (23). Details of the problem transformation are described below, which consists of four steps.
(a) Transformation of Individual Fairness Constraints. Before employing Lyapunov optimization, we remove the absolute value sign from the constraint in Eq. (9) and rephrase it as two equivalent constraints,
@@ III-C Transformation Under Lyapunov Optimization
(b) Introduction of Virtual Queues $Z_{i}(t)$, $Q_{i}(t)$ and $\Theta(t)$. Following the general process of Lyapunov optimization [35], we define two virtual queues, namely $Z_{i}(t)$ for the constraint in Eq. (11) and $Q_{i}(t)$ for the constraint in Eq. (12), for all clients $i\in\mathbb{N}$. Specifically, these queues are initialized as $Z_{i}(0)=0$ and $Q_{i}(0)=0$, and updated according to the following rule
@@
By the introduction of $Z_{i}(t)$ and $Q_{i}(t)$, we have Theorem 2, which converts the long-term constraints on $x_{i,t}$ in Eq. (11) and (12) into the stability constraints for $Z_{i}(t)$ and $Q_{i}(t)$, respectively. The proof of Theorem 2 is provided in the supplementary file.
@@
Then, we define a global queue $\Theta(t)$, which stores the state of $Z_{i}(t)$ and $Q_{i}(t)$ for all clients. That is,
@@
(c) Introduction of Lyapunov Function $L\left(\Theta(t)\right)$ and Lyapunov Drift $\Delta\left(\Theta(t)\right)$. Following the general process of Lyapunov optimization [35], we define the Lyapunov function [35] as
@@
(d) Introduction of the Tradeoff Factor $V$ and two Variables $m_{i,t}$ and $n_{i,t}$. To ensure $\Delta\left(\Theta(t)\right)$ remains sufficiently small, a straightforward method is to combine the optimization objection with the drift $\Delta\left(\Theta(t)\right)$, and minimize both simultaneously. Mathematically, in our work, the optimization problem in Eq. (10) is transformed into
@@
The Final Formulation. Based on Theorem 3, the problem P2 in Eq. (18) is transformed into
@@
which is the final optimization function. In the following, we select clients by minimizing $G(\mathbb{S}^{t})$ in Eq. (23) in each round.
@@ IV-A The Client Selection Strategy
5:     Determine the reference client $i^{\star}$ for $i\in\mathbb{P}_{k}$
@@
Specifically, a set function $g:2^{\mathbb{N}}\rightarrow\mathbb{R}$ is submodular if for every $A\subseteq B\subseteq\mathbb{N}$ and $\𝐵i\in\mathbb{N}\backslash B$ it holds that $g(A\cup\{i\})-g(A)>g(B\cup\{i\})-g(B)$. One typical example of submodular function is the facility location function [37]. Suppose we aim to select locations from a set of positions =$\mathbb{N}=\{1,\ldots,N\}$ to open facilities and to serve a collection of $K$ users. If a facility is located at position $j$, the service it provides to user $i$ is quantified by $M_{i,j}$. Each user is assumed to select the facility with the highest service, and the total service provided to all users is modeled by the set function
$f(\mathbb{S})=\sum_{i=1}^{m}\max_{j\in\mathbb{S}}M_{i,j},$ (24)
$\overline{G}(\mathbb{S}^{t})=G(\{e\})-G(\mathbb{S}^{t}\cup\{e\}),$ (25)
@@ IV-B Convergence Analysis
To implement the theoretical analysis, we establish six assumptions regarding the local models and data distribution heterogeneity among clients. The analysis uses FedAvg as the aggregation method, and it can be extended to other federated optimization methods as well.
First, we assume that the estimation error between the client subset and the full client set (Eq. (2)) is small and can be quantified by a variable $\rho$, as stated in Assumption 1. Note that $\rho$ is used as a measure to characterize the quality of the estimation, and the analysis holds for any $\rho<\infty$.
Assumption 1. At a round $t$, we assume that the gradient aggregated from the selected subset of clients can provide a good approximation of the gradient aggregated from the full set, i.e.,
@@
Based on these assumptions, the proposed client selection strategy is demonstrated to converge to the global optimal parameter $\bm{w}^{\star}$ at a rate of $\mathcal{O}(1/t)$ for heterogeneous data settings, as stated in Theorem 4. The proof is provided in the supplementary file. The convergence rate $\mathcal{O}(1/t)$ is the same as loss-guided client selection methods [16, 12, 40].
@@
In Eq. (2), the term $\rho$ encodes the estimation error. When more clients are selected, the term $\rho$ is smaller. Particularly, $\rho$ becomes zero when all clients are selected, i.e., $K=N$. In practical settings, as only limited clients can be selected, $\rho$ remains a non-vanishing term. In our experiments, we also observe that there exists a non-diminishing solution bias dependent on $\rho$. This observation is consistent with our theoretical analysis. The impact of varying $K$ is also empirically analyzed in [41].
@@ V-A Experimental Settings
We consider the cross-device federated learning scenario where $N=100$ clients exist, each with limited computational power. The FedAvg method [1] is used as the aggregation method. We evaluate our method on two datasets, FMNIST [42] and CIFAR-10 [43]. Following [12], we opt for basic and small-scale models to accommodate the clients’ restricted computational resources. Specifically, for the FMNIST dataset, we utilize a multilayer perception (MLP) with two hidden layers. For CIFAR-10, we adopts a convolutional neural network (CNN) architecture consisting of three convolutional layers. Details of training and hyperparameter settings can be found in supplementary file.
@@
Data Partition Methods. We explore four data partitioning methods. The first is an independently and identically distributed (IID) approach, where we randomly partition the dataset into $N$ parts, with each client assigned one part. We also consider three heterogeneous data partitioning approaches as follows.
@@
(i) 1 Shard Per Client (1SPC). Following [1], we divide the dataset into $N$ shards, ensuring that data within a shard shares the same label. We randomly assign a shard to each client so that a client has data with only one label. In this case, we select $K=10$ clients in each round.
@@
(ii) 2 Shards Per Client (2SPC). The dataset is partitioned into $2N$ shards, with each shard containing data that shares the same label. Clients are randomly assigned two shards, allowing them to have data with as most two distinct labels. In this case, we set $K=5$ in each round.
@@
(iii) Dirichlet Distribution (Dir). We partition the dataset based on a Dirichlet distribution parameterized by a concentration variable $\alpha$, where a smaller value of $\alpha$ indicates higher heterogeneity. In our work, we set $\alpha=0.8$, and determine the data size of each label for each client following [44, 12]. In this case, we choose $K=5$ clients in each round. We also experimented with smaller values, such as $\alpha=0.5$ and $\alpha=0.1$. We found that our method consistently achieved the best model performance.
@@
Note that the 1SPC and 2SPC scenarios address label shift heterogeneity, while the Dir scenario accounts for heterogeneity in both labels and dataset size.
Figure 2: Test accuracy under the IID scenario.
Comparison Methods. We first evaluate the effectiveness of the proposed method by comparing it with existing client selection strategies in terms of model performance. The baselines are as follows. (1) Random selection strategy (Random, 2017) [1]; (2) Loss-guided selection methods, including active client selection strategy (AFL, 2019) [10], power-of-choice selection strategy (Powerd, 2022) [11], and diverse client selection strategy (divFL, 2022) [16]; (3) Contribution-based methods, including the Shapley value-based method (ShapleyFL, 2023) [22] and ranking-based client selection (FedRank, 2024) [45]. Cluster-based methods are not considered, as they require explicit clustering patterns, which are not applicable in the 2SPC and Dir scenarios. Additionally, we include full participation method (Full) with $K=100$ as a reference.
Furthermore, we evaluate the proposed method in terms of fairness. As introduced in Section II, the uniform selection constraint from [29] (2023) in performance fairness research is most aligned with our work. In our study, we adopt this uniform selection constraint and apply it to loss-guided selection methods, given their superior model performance. Specifically, we apply the uniform selection constraint to AFL [10], Powerd [11], and divFL [16], denoting them as AFL+Fair, Powerd+Fair, and divFL+Fair, respectively. We then compare our method against these three baselines in terms of fairness.
@@
Note that the divFL [16] method also selects clients to best represent full client participation, which aligns with our proposed Principle I. However, divFL does not consider fairness. By comparing divFL with our proposed longFed, we can evaluate the impact of our proposed individual fairness on improving model performance. Additionally, by comparing divFL+Fair with longFed, we can further evaluate how our proposed individual fairness outperforms the uniform selection constraint.
@@ V-B Experimental Results
V-B1 Model Performance
The experimental results under the IID scenario are presented in Fig. 2. First, as shown in Fig. 2 (a) and (b), our proposed longFed outperforms existing methods, particularly on CIFAR-10 dataset, validating its effectiveness in enhancing model performance in the IID scenario. Second, longFed consistently outperforms divFL, further demonstrating the effectiveness of our proposed individual fairness in the IID scenario.
Figure 4: Fairness results on the CIFAR-10 dataset in the 2SPC scenario.
The experimental results under three heterogeneous scenarios are presented in Fig. 3. First, compared to prior works (except for divFL), the proposed longFed exhibits faster convergence and superior test accuracy. Notably, it achieves an approximate $20\%$ improvement in the 1SPC scenario and an $8\%$ improvement in the 2SPC and Dir scenarios on the FMNIST dataset. Besides, the proposed method consistently achieves performance comparable to full client participation across all three scenarios and both datasets. This validates the effectiveness of our method in enhancing model performance. Furthermore, compared to divFL, longFed exhibits similar performance in Fig. 3 (a), (c), and (d), but achieves significant improvements in the other three scenarios. This highlights the effectiveness of our proposed individual fairness in improving model performance.
Notably, in Fig. 3 (b) and (f), longFed even surpasses full client participation in some rounds. This phenomenon typically occurs in the early stages of training due to optimization dynamics and the stochasticity introduced by client selection. As training progresses, these effects gradually diminish. Additionally, FedRank performs the worst, as it relies solely on pairwise relationships between clients and lacks a global perspective on data diversity. This leads to biased selections-for example, it tends to overlook important yet less frequently ranked clients-ultimately resulting in suboptimal training and poorer convergence.
V-B2 Fairness